OpenStudy (anonymous):

What number doesn't belong in this set...name the other 4 numbers that are the same

OpenStudy (anonymous):

1/4 -7 SQUARE ROOT OF 25 8/2 and cube root of 1

OpenStudy (anonymous):

....

OpenStudy (anonymous):

you are not getting an answer because it could be literally anything what does it mean "what number does not belong in this set"?

OpenStudy (anonymous):

you have \(\frac{1}{4},-7,5,4,1\)

OpenStudy (anonymous):

like the types of number....irrational...whole...natural..rational etc

OpenStudy (anonymous):

one of those numbers is a fraction one is negative who knows?

OpenStudy (anonymous):

ok but i have another question....

OpenStudy (anonymous):

is 0.6 repeating a rational or irrational number

OpenStudy (anonymous):

maybe the negative one doesn't belong maybe the fraction doesn't belong the question is not well posed

OpenStudy (anonymous):

it is rational .6 repeating is \(\frac{6}{9}=\frac{2}{3}\)

OpenStudy (anonymous):

any repeating decimal is rational

OpenStudy (anonymous):

but what if it is .45678 repeating

OpenStudy (anonymous):

then that is rational as well

OpenStudy (anonymous):

as long as it repeats, it is rational

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

HOWEVER if you mean \[.456789101112131415...\] then this is not repeating it just has a pattern

OpenStudy (anonymous):

but if you mean \[.456784567845678...\] then yes, it is rational

OpenStudy (dape):

Well all are integers except \(\frac{1}{4}\).

OpenStudy (dape):

But I agree with satellite73, it's a vague question. All are positive real numbers except -7.